OpenStudy (anonymous):

what is the line that passes through (4,3) and that has a slope of 2 ?

OpenStudy (anonymous):

???????????

OpenStudy (anonymous):

y=mx+b slope or m is given which is 2 x= 4 y= 3 therefore, \[3=(2)4+b\] \[3=8+b\] b=-5

OpenStudy (anonymous):

y=2x-5